OpenStudy (anonymous):

Angle A and Angle B are coplementary angles. If angle A is six more than three times angle B, what are the measures of the angles?

OpenStudy (ash2326):

@bev199 Do you know the meaning of complimentary?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

complementary angles are those whose sum is 90 degrees.Use this two write a linear equation in one variable and solve.

OpenStudy (ash2326):

Two angles are complimentary if their sum is 90 degrees

OpenStudy (ash2326):

@bev199 try to create the equation

OpenStudy (anonymous):

A+b=90

OpenStudy (ash2326):

you're right, now create the second equation

OpenStudy (anonymous):

3b+6=A

OpenStudy (ash2326):

Yeah you're right. Now can you find A and B from both these equations

OpenStudy (anonymous):

So would it be 3(45)+6(45)=A

OpenStudy (anonymous):

no.i think you got it wrong.It is A + B=90. now substitute A=3B+6 in the equation and solve to find A.

OpenStudy (ash2326):

We have two equations \[A+B=90\] \[A=3B+6\] Now substitute A from second to the first equation \[3B+6+B=90\] Find B from this and then find A:) Can you do it now?

OpenStudy (precal):